OpenStudy (anonymous):

3. find the exac value of each follwing. do not use a calculator A. cos(7pi/12) part1: find two common angles that either add up to or differ to 7pi/12. rewite this problem as the cosine of either a sum or a difference of those two angles Part2: evaluate the expression

OpenStudy (anonymous):

use identity cos(A+B)= cos(a)cos(b)-Sin(a)sin(b) 7pi/12 =(pi/4+pi/3)

OpenStudy (anonymous):

(rad2-6)/4